the discounted ticket price for morning movie is $5 the average cost of a full price movie ticket $11 a total of forty-one people bought tickets to the movies on Monday and the theatre earned a total of $307 how many discounted and full price tickets did the theatre sell?
Answer by addingup(3677)

d + f = 41 therefore f = 41-d Use this value for f next:
5d + 11f = 307
5d + 11(41-d) = 307
5d + 451 - 11d = 307
-6d = -144 divide both sides by -6 and remember that -/- = +
d = 24
they sold 24 discount tickets
and
41 - 24 = 17 full price tickets
.
Happy learning